OpenStudy (anonymous):

Find the exact value of sin(a+b): If sina=3/5, 0< a< pi/2; cos B= 2 root 5/5, =pi/2

OpenStudy (butterflydreamer):

hmm.. sin ( a+ b) = sina cosb + sinb cosa We are given: sina = 3/5 , cos b = 2 root 5/5 So draw out two triangles: